OpenStudy (anonymous):

Differential Equation y'-y=2x-3 solve with as a first order equation

OpenStudy (anonymous):

The final answer is 2x+y-1=ce^x

OpenStudy (anonymous):

Any thoughts fellas?

OpenStudy (anonymous):

@Godlovesme ?

OpenStudy (anonymous):

@Justin_Lee ?

OpenStudy (justin_lee):

Wasn't what I got. How did you get that answer?

OpenStudy (anonymous):

from the book

OpenStudy (justin_lee):

Sorry. You are right. I copied down the original equation wrong when I was solving it. I thought it was y'+y.

OpenStudy (anonymous):

Did you actually solve it?!

OpenStudy (anonymous):

@Justin_Lee

OpenStudy (justin_lee):

Yeah. Have you taken differential equations before?

OpenStudy (anonymous):

not at all

OpenStudy (justin_lee):

What level of math are you?

OpenStudy (anonymous):

tbh it was only our second lesson

OpenStudy (justin_lee):

Okay. Do you know the method of integrating factors?

OpenStudy (justin_lee):

Sorry. I haven't eaten in a while, so I was getting food.

OpenStudy (anonymous):

sort of, and first for it, boun appetite

OpenStudy (anonymous):

with all due resfect, it's more important

OpenStudy (anonymous):

respect

OpenStudy (justin_lee):

Thanks. Okay. This can be solved using either the method of integrating factors or using a characteristic equation.

OpenStudy (justin_lee):

The method of integrating factor requires you to multiply the entire equation by a dummy expression, say u(x).

OpenStudy (justin_lee):

This gives you: y'(x)*u(x) - y(x)*u(x) = 2x*u(x) - 3u(x).

OpenStudy (justin_lee):

This allows you to take advantage of the product rule for the left hand side of the equation. Remember, the product rule is: (fg)' = g'f + f'g.

OpenStudy (justin_lee):

That now kind of resembles the left hand side of the expression: y'(x)*u(x) - y(x)*u(x). This assumes that u(x) = f and y(x) = g. Now, notice something interesting here. u'(x) = u(x), if we are following the pattern set by the product rule.

OpenStudy (justin_lee):

You can recognize that only one function can do this: e^x. Or, you can prove this to yourself in the following manner: u'(x) = du(x)/dx = u(x) du(x)/u(x) = dx Integrating both sides gives: ln(u(x)) = x (the constant is irrelevant) Raising both sides by an exponential yields: u(x) = e^x.

OpenStudy (justin_lee):

Now, you can simplify the left hand side from: y'(x)*u(x) - y(x)*u(x) to: (y(x)*u(x))' by following the steps backwards from the product rule: g'f + f'g = (fg)' Keep in mind that u(x) = e^x.

OpenStudy (justin_lee):

Does all this make sense so far? Also, I made a mistake in the second to last post. Can you spot it?

OpenStudy (anonymous):

no, can you tell me where is it? sorry it's too late here to think that hard lol

OpenStudy (anonymous):

but @Justin_Lee I'll look at that in the morning thank you so much!

OpenStudy (anonymous):

such a vivid explanation!

OpenStudy (justin_lee):

u(x) should be e^-x.

OpenStudy (justin_lee):

Sure. I will type up some concluding remarks and let you be on your way.

OpenStudy (anonymous):

wow! Truly thanks a lot!!!!!!

OpenStudy (anonymous):

It has been a long day

OpenStudy (justin_lee):

u(x) should be e^-x because of the y' - y. Imagine that u(x) = e^ax. Then, if we multiply some y(x) by u(x) and take the derivative of that product, we would yield: y'(x)*u(x) + a*y(x)*u(x) [prove it to yourself]. Now, having simplified the expression by reversing the product rule, we have: (y(x)*u(x))' = 2x*u(x) - 3*u(x). Replace u(x) with e^-x gives: (y(x)*e^-x)' = 2x*e^-x - 3*e^-x Integrate both sides by x to get rid of the derivative on the left hand side (don't forget to tack on a constant to the right hand side). This gives: y(x)*e^-x = -2x*e^-x + e^-x + c [solve it out to prove it to yourself that this is the case] Divide by e^-x in both sides to get the answer that you found in the book.

OpenStudy (justin_lee):

Alternatively, you can use the DE's characteristic equation in conjunction with the method of undetermined coefficients (which is what people usually use to solve these types of problems). This method sometimes requires a leap of faith at first in order to believe your result, but is much faster than the other method I showed you. If you are working with multiple derivatives, expect to learn this in the near future.